OpenStudy (anonymous):

A circle is centered at (3, 8) and is tangent to the x-axis. Write its equation in standard form.

OpenStudy (serenity74):

My first question what is the equation of a circle?

OpenStudy (anonymous):

it doesn't say this is all it gives me

OpenStudy (serenity74):

Okay, you use (x-h)^2 +(y- k)^2 = r^2 given points (3,8)

OpenStudy (serenity74):

Wait is this Calculus?????

OpenStudy (anonymous):

thanks and geometry

OpenStudy (serenity74):

Oh gosh you scared me for a second lol and yea you just use that formula and do you know how to solve for the Radius??

OpenStudy (anonymous):

they put me in the wrong class so i dont know what im doing and no

OpenStudy (serenity74):

Lol I hate that! Happened to me with pre-calc but thankfully I got to stay in pre-calc. Did they not give you the radius? Cause usually they give you the radius in the question.

OpenStudy (anonymous):

they placed me in acc geometry and i said cp geometry and then when they gave me my schedule they said they couldnt change me .-. and no like the question i wrote is the exact same thing of what they gave me

OpenStudy (serenity74):

Hmmm let me review my notes real quick okay? I will get you the answer I swear or this will drive me nuts too lol.

OpenStudy (anonymous):

ok lol thanks

OpenStudy (serenity74):

I think after you plug in the coordinates into the equation you distribute it and then equal it to 0. After you are done with that let me know and show me the work you have done so far so i can see whether what you did was right or not.

OpenStudy (anonymous):

so do you mean like foil it?

OpenStudy (serenity74):

Yup!

OpenStudy (anonymous):

so far i have x^2 - 6x + y^2 - 16y + 73 = 0

OpenStudy (serenity74):

Yup, now do you know how to do completing the square method? I would explain it to you after you read the link i am about to send you. Just so you can get a glance off it of it mathematically and then my gibberish XD

OpenStudy (anonymous):

yeah i know im doing it right now lol and thanks so much i really appreciate it

OpenStudy (serenity74):

No problem! Let me know if you need anything us in the future! I will try my best in helping you :) Have a great day and peace.